# Asymptotic Behavior of Rupture Solutions for the Elliptic MEMS Equation with Hénon-Type Term

### v1

@@
Abstract: For an elliptic MEMS equation with Hénon-type $\Delta u=\dfrac{\lambda|x|^{\alpha}}{u^{p}}+F$. we study rupture solutions (i.e. solutions for which $u(x_{0})=0$ at some point $x_{0}$, also $x_{0}$ is called a rupture point). In this paper we focus on the special case where the rupture occurs at the origin. According to the different Hénon-type exponents $\alpha$, we analyze the asymptotic behavior of such solutions near the origin, derive a full asymptotic expansion of arbitrary order in a neighborhood of the origin. Moreover, with respect to radial solution and non-radial solution with asymptotic radial condition, we prove both of them exist near the rupture point by constrcuting it.
Keywords:Asymptotic expansions, Rupture point, Hénon-type.
MSC:
@@ 1 Introduction
Micro-electro-mechanical systems (MEMS) are devices that integrate miniature components by combining electrostatic effects with microfabrication technology. Nowadays MEMS devices are widely used in electronic equipment, aerospace engineering and medical applications. With the rapid development of high-tech fields such as artificial intelligence and ultra-precision machining, the performance requirements on MEMS have been increasing; however, electrostatic MEMS typically suffer from pull-in instability.
Taking the electrostatically actuated elastic membrane system considered in this paper as a prototype, we proceed as follows. As illustrated in Figure 1, a deformable elastic membrane is put above a fixed ground plate. When a voltage is applied, the membrane is attracted towards the plate. As the voltage increases, the membrane eventually touches the plate. This so-called pull-in (or touchdown) instability may affect the stable operation of high-precision microelectronic devices and may even cause irreversible damage. On the other hand, in some applications such contact is actually desired. For instance, the triggering of air bag restraint system or the operation of printers. In this work we focus on the profile of the membrane near the touchdown point at the moment when contact occurs.
@@
We consider the following model. Let $u$ denote the distance between the membrane and the ground plate, $\lambda$ the applied voltage, and $|x|^{\alpha}$ the dielectric profile on the membrane (representing the spatial inhomogeneity induced by the variation of the dielectric constant). We are interested in the situation where the membrane touches the ground plate at $x=0$. In this case the governing main equation can be written as $\Delta u=\dfrac{\lambda|x|^{\alpha}}{u^{2}},x\in\mathbb{R}^{2}\setminus\{0\}$ ( $\lambda>0$ and $\alpha>-2$). In this paper we are interested in a more general equation. More precisely, we generalize the negative exponent $p=2$ in the right-hand side of the equation to a general $p>0$ with $p>1/2\alpha$, introduce an additional force term $F$, and extend the above two-dimensional model to the NN-dimensional setting($N\geq 1$). In the present work we focus on the following problem:
$\begin{cases}\Delta u=\dfrac{\lambda|x|^{\alpha}}{u^{p}}+F,&x\in\mathbb{R}^{N}\setminus\{0\},\\[4.0pt] u(0)=0,&x\in\mathbb{R}^{N}\setminus\{0\},\\ u\geq 0,&x\in\mathbb{R}^{N}\setminus\{0\},\end{cases}$ (1)
(where $\lambda>0$ , $2<\alpha<2p$ , $p>0$ are constants).
We mainly focus on the existence of asymptotically radially symmetric solutions and their asymptotic behavior near the rupture point with the influence of the Hénon-type term (the $|x|^{\alpha}$ term) of (1).
(Definition of asymptotic radial rupture solution: let $u(x)=u(r,\theta)$ be a nonnegative rupture solution of (1). If there exists $\beta$ such that $\lim_{r\to 0^{+}}r^{\beta}u(r,\theta)=f$ in $C^{2}(S^{N-1})$ for some constant $f$, then we say that $u(x)$ is asymptotic radial rupture solution near the origin.)
For the equation $\Delta u=u^{-p}$ with $p>0$ in $\mathbb{R}^{N}$ $(N>2)$, it has been proved in [6] that for any given $a>0$ there exists a unique radially symmetric solution $u=u(r)$ satisfying $u(0)=a$ and $u(r)\to\infty$ as $r\to\infty$.For the specfic case of $N=2$ and $F=0$ in (1).[11] investigated the isotropic and anisotropic behaviors of rupture solutions in different parameter regions, also the author studied global solutions on $R^{2}\setminus{0}$. For specfic case of $N=2$ in (1), and for different values of $\alpha$, [8] used phase-plane analysis together with the Lojasiewicz–Simon inequality to classify the asymptotic behavior of rupture solutions at the origin in the cases of asymptotically isotropic and asymptotically anisotropic profiles. Moreover, for solutions that are asymptotically isotropic at the origin, [8] obtained a more precise description and derived the first two terms in the asymptotic expansion of rupture solutions. For the specfic case $F=0$ in (1), [12]told that $u(|x|)\equiv\left(\frac{1}{\frac{\alpha+2}{p+1}\left(\frac{\alpha+2}{p+1}+N-2\right)}\right)^{\frac{1}{p+1}}.$ in radial cases.
For the asymptotic expansion of solutions near the fixed point. [4] use spheric harmonics to derive the asymptotic expansion near an isolated singularity for the Yamabe equation $\Delta u=\tfrac{1}{4}n(n-2)u^{\frac{n+2}{n-2}}.$ More recently, [9] also use spheric harmonics to analyze multi-term asymptotic expansions of the steady thin-film-type equation $\Delta u=u^{-p}-q,(x\in B_{R}\setminus\{0\}\quad p,q\in\mathbb{R},\ p>0)$ near the rupture point (the origin).
Although there have been some results on the existence of rupture solutions to MEMS-type elliptic equations and on their asymptotic behavior near the rupture point, to the best of our knowledge no corresponding results are available for the equation with the Hénon-type term considered here. In this paper we perform a detailed case-by-case analysis and establish the theoretical framework and results in this direction, which yields richer results. Now we give the main thereom
Theorem $1.$
For the elliptic MEMS equation (1) with $N\geq 2$, there exists at least one radial solution near the origin satisfying
$u(r)=\Lambda r^{\frac{\alpha+2}{p+1}}+\sum_{i=1}^{\infty}d_{i}r^{\frac{(2p-\alpha)i+(\alpha+2)}{p+1}}=u_{s}(r)\left(1+O\left(r^{\frac{2p-\alpha}{p+1}}\right)\right)\qquad\text{as }r\rightarrow 0^{+},$
where $\mathrm{u}_{s}(r)=\Lambda r^{\frac{\alpha+2}{p+1}}$ and $\Lambda=\left(\frac{\lambda}{\frac{\alpha+2}{p+1}\left(\frac{\alpha+2}{p+1}+N-2\right)}\right)^{\frac{1}{p+1}}.$
Remark. For the case of $N=1$, there exists at least one solution near the origin satisfying
$u(x)=\Lambda r^{\frac{\alpha+2}{p+1}}+\sum_{i=1}^{\infty}d_{i}x^{\frac{(2p-\alpha)i+(\alpha+2)}{p+1}}=u_{s}(x)\left(1+O\left(x^{\frac{2p-\alpha}{p+1}}\right)\right)\qquad\text{as }x\rightarrow 0^{+},$
where $\mathrm{u}_{s}(x)=\Lambda r^{\frac{\alpha+2}{p+1}}$ and $\Lambda=\left(\frac{\lambda}{\frac{\alpha+2}{p+1}\left(\frac{\alpha+2}{p+1}-1\right)}\right)^{\frac{1}{p+1}}.$, since the arguments for radial solution in the case $N\geq 2$ is completely similar to these for $N=1$, we shall omit the case $N=1$.
@@
For the elliptic MEMS equation (1), there exist infinitely many non-radial positive asymptotic radial rupture solutions satisfying
$u(x)=\Lambda|x|^{\frac{\alpha+2}{p+1}}+C_{10}\left(\frac{x}{|x|}\right)|x|^{u_{1}}+\sum_{j=2}^{\infty}\sum_{i=0}^{i-1}C_{ji}\left(\frac{x}{|x|}\right)(\ln|x|)^{i}|x|^{u_{j}}=u_{s}(|x|)\left(1+O\left(|x|^{u_{1}}\right)\right)$
as $|x|\rightarrow 0^{+}$,
here $\mathrm{u}_{s}(|x|)=\Lambda|x|^{\frac{\alpha+2}{p+1}}$, $\Lambda$ is the same as above, $\left\{\mu_{j}\right\}_{j\geq 1}$ is a strictly increasing sequence of positive numbers diverging to $\infty$. Where
[…]
in the case $F\neq 0$.
[…]
@@
And $\delta^{(k)}=\frac{-1}{2}(p+1)(N+2)+\frac{1}{2}\sqrt{(N-2)^{2}(p+1)^{2}+4(p+1)k(N-2+k)},$
@@
Let $t=\ln|x|$ $\theta=\dfrac{x}{|x|}$, then
$\Delta u=u_{rr}+\frac{N-1}{r}u_{r}+\frac{1}{r^{2}}\Delta_{\theta}u=\frac{\lambda|x|^{\alpha}}{u^{p}}+F$ (2)
Let $\dfrac{\alpha+2}{p+1}=q$, $z(t,\theta)=r^{-\frac{\alpha+2}{p+1}}u(x)-\Lambda$,    where $\Lambda=\left(\frac{\lambda}{\frac{\alpha+2}{p+1}\left(\frac{\alpha+2}{p+1}+N-2\right)}\right)^{\frac{1}{p+1}}.$. then we can obtain:
$z_{tt}+\left(N-2+2q\right)z_{t}+(\alpha+2)(N-2+q)z+\Delta{}_{\theta}z=\frac{\lambda}{(z+\Lambda)^{p}}-\frac{\lambda}{\Lambda^{p}}+\frac{\lambda pz}{\Lambda^{p+1}}+Fr^{pq-\alpha}$
Set $f(z)=\dfrac{\lambda}{(z+\Lambda)^{p}}-\dfrac{\lambda}{\Lambda^{p}}+\dfrac{\lambda pz}{\Lambda^{p+1}}$, so
$z_{tt}+\left(N-2+2q\right)z_{t}+(\alpha+2)(N-2+q)z+\Delta{}_{\theta}z=f(z)+Fe^{\frac{2p-\alpha}{p+1}t}$ (3)
@@
Suppose that $u(x)$ is a radially symmetric solution of (1) with $F\neq 0$ and $\alpha<2p$. Assume it exists $\varepsilon>0$ such that
$u(r)=u_{s}(r)\left(1+O\left(r^{\varepsilon}\right)\right),~~as~r\rightarrow 0^{+},$
where $\mathrm{u}_{s}(r)=\Lambda r^{\frac{\alpha+2}{p+1}}$ and
$\Lambda=\left[\frac{\lambda(\alpha+2)}{p+1}\left(N-2+\frac{\alpha+2}{p+1}\right)\right]^{\frac{1}{p+1}}.$
Set
$z(t)=r^{-\frac{\alpha+2}{p+1}}u(r)-\Lambda,\quad t=\ln r.$
@@
Proof. First, we have
@@
$f(z)=\lambda(\Lambda+z)^{-p}-\lambda\Lambda^{-p}+\lambda pz\Lambda^{-(p+1)}=O(z^{2})\quad~~as~z(t)\rightarrow 0$
@@
case1. $\alpha>\frac{(p+1)(N-2)}{2}(\sqrt{1+\frac{1}{p}}-1)-2$
$\left\{\begin{aligned} \sigma_{1}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)+\frac{i}{2}\sqrt{4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)-\left(N-2+2\frac{a+2}{p+1}\right)^{2}},\\ \sigma_{2}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)-\frac{i}{2}\sqrt{4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)-\left(N-2+2\frac{a+2}{p+1}\right)^{2}},\end{aligned}\right$.
case2. $\alpha=\frac{(p+1)(N-2)}{2}(\sqrt{1+\frac{1}{p}}-1)-2$
@@
case3. $2<\alpha<\frac{(p+1)(N-2)}{2}(\sqrt{1+\frac{1}{p}}-1)-2$
$\left\{\begin{aligned} \sigma_{1}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)+\frac{1}{2}\sqrt{\left(N-2+2\frac{a+2}{p+1}\right)^{2}-4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)}<0,\\ \sigma_{2}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)-\frac{1}{2}\sqrt{\left(N-2+2\frac{a+2}{p+1}\right)^{2}-4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)}<0,\end{aligned}\right$.
@@
$z(t)=O\left(e^{\frac{2p-\alpha}{p+1}t}\right)\qquad\text{as }t\to$-$\infty$
@@
$z(t)=B_{1}e^{\sigma_{1}^{(0)}}t\int_{-\infty}^{t}e^{-\sigma_{1}^{(0)}s}\left[f(z(s))+Fe^{\frac{2p-\alpha}{p+1}t}\right]ds+B_{2}e^{\sigma_{2}^{(0)}}t\int_{-\infty}^{t}e^{-\sigma_{2}^{(0)}s}\left[f(z(s))+Fe^{\frac{2p-\alpha}{p+1}t}\right]ds$ (4)
[…]
@@
[…]
@@
Therefore the prove of Theorem2.1 is completed.

### v3

# Asymptotic Behavior of Rupture Solutions for the Elliptic MEMS Equation with Hénon and External Pressure Terms

@@
Abstract: This paper investigates an elliptic MEMS-Type equation with Hénon and external pressure terms:
[…]
where $N\geq 1,\lambda>0,p>0,\alpha>-2$ and $F\in R$ are constants. We study positive rupture solutions with rupture point at the origin ($u(0)=0$). Our main emphasis is on asymptotic radial rupture solutions: we prove the existence of both radial and non-radial solutions, characterize their asymptotic behavior near the origin, and obtain a full asymptotic expansion of arbitrary order
Keywords: Asymptotic expansions, MEMS-Type Equation, Hénon term.
MSC2020: 35J75, 35J61, 35C20, 35B40, 74K15, 74G70.
@@ 1 Introduction
In this paper, we investigate the following elliptic MEMS-type equation with Hénon term:
[…]
where $N\geq 1,\lambda>0,p>0,\alpha>-2$ and $F\in R$ are constants. We establish the existence of solutions to (1) and derive asymptotic expansions in a sufficiently small neighborhood of the origin.
Micro-Electro-Mechanical Systems (MEMS) are devices that integrate miniature mechanical components with electronics using microfabrication technology. We consider a canonical MEMS model, as illustrated in Figure 1 (c.f.[24]), consisting of a deformable elastic membrane suspended above a fixed ground plate. When a voltage is applied, the electrostatic force attracts the membrane towards the plate. As the voltage increases, the membrane may eventually touch the plate. This phenomenon, known as ”pull-in instability”, can compromise the stable operation of high-precision microelectronic devices. Conversely, in certain applications such as airbag restraint systems or inkjet printers, such contact is desirable.
In this work, we focus on the profile of the membrane near the touchdown point—referred to mathematically as the rupture point—at the moment contact occurs. A systematic characterization of the local asymptotic profile near the rupture point, along with its dependence on system parameters, provides essential guidance for the design and reliability assessment of MEMS devices.
@@
Mathematically, let $u\geq 0$ denote the distance between the membrane and the ground plate. Let $\lambda>0$ represent the applied voltage, $|x|^{\alpha}~(\alpha>-2)$ represent the variable permittivity profile of the membrane, and $F$ denote the external pressure (where $F>0$ corresponds to a downward force and $F<0$ to an upward force). We focus on the case where the membrane ruptures at $x=0$. This is modeled by the equation (cf. [23, 24, 25]):
[…]
Problem (1) is obtained by generalizing the exponent $p=2$ to a general $p>0$, extending the domain to the N-dimensional setting ($N\geq 1$).
Problem (1) has been studied in various parameter regimes. For $\alpha=0$, the rupture solutions of problem (1) were investigated in [14, 11]. In particular, for $\lambda=1$, $p>1$ and $N\geq 2$, the Hölder continuity and the Hausdorff dimensions of rupture sets were studied in [11]. For $\lambda=1$, $F=0$ and $N\geq 3$, [14] obtained infinitely many non-radial rupture solutions for $p>\frac{N-1-2\sqrt{N-2}}{2\sqrt{N-2}-(N-5)}$. For $\alpha\neq 0$, the rupture solutions of problem (1) were investigated in [8, 12, 13]. In particular, for $N=2$, $p=2$ and $F\geq 0$, [8] investigated the isotropic and anisotropic singularities of rupture solutions with respect to different values of $\alpha$. For $N=2$ and $F=0$, [12] analyzed the isotropic and anisotropic behaviors in various parameter regions of $(\alpha,p)$ and studied global solutions on $\mathbb{R}^{2}\setminus\{0\}$. For $\lambda=1$ and $F=0$, a dichotomy result regarding the Morse index of such solutions was presented in [13], depending on whether the exponent $p$ is below or above the critical threshold .
Significant progress has also been made regarding the asymptotic expansion of solutions near singular points. For instance, [8] derived the first two terms of the expansion for asymptotically isotropic rupture solutions. Using spherical harmonics, [4] derived the asymptotic expansion near an isolated singularity for the Yamabe equation $\Delta u=\frac{1}{4}n(n-2)u^{\frac{n+2}{n-2}}$. More recently, [9] utilized spherical harmonics to analyze asymptotic expansions to arbitrary order for the steady thin-film-type equation near the origin. This corresponds to a special case of Equation (1) with $\lambda=1$ and $\alpha=0$.
In this paper, we generalize these analyses to the general $\alpha$. Regarding asymptotic radial solutions to problem (1), we investigate the existence and asymptotic behavior of radial and non-radial rupture solutions near the origin. An asymptotic radial solution is defined as a solution of the form
$u(x)=C|x|^{d}\left(1+O\left(|x|^{\epsilon}\right)\right)\quad as~|x|\rightarrow 0^{+}$
for some $\epsilon>0$, where $C$ and $d$ are non-zero constants. The main results are stated as follows.
@@
For the elliptic MEMS problem (1) with $N\geq 2$, $F\neq 0$ and $2<\alpha<2p$, there exists at least one radial rupture solution near the origin satisfying
$u(r)=\Lambda r^{\frac{\alpha+2}{p+1}}+\sum_{i=1}^{\infty}d_{i}r^{\frac{(2p-\alpha)i+(\alpha+2)}{p+1}}=\Lambda r^{\frac{\alpha+2}{p+1}}\left(1+O\left(r^{\frac{2p-\alpha}{p+1}}\right)\right)\qquad\text{as }r\rightarrow 0^{+},$
where $d_{1}$ depends on $F,p,\alpha,N$ and $d_{2},\dots,d_{\infty}$ depend on $p,\alpha,N,\lambda$.
$\Lambda=\left(\frac{\lambda}{\frac{\alpha+2}{p+1}\left(\frac{\alpha+2}{p+1}+N-2\right)}\right)^{\frac{1}{p+1}}.$ (2)
Remark.
For the case of $F=0$, it has been proved in [13] that $u(x)=\Lambda|x|^{\frac{\alpha+2}{p+1}}$ is the only positive radial rupture solution of (1).
Theorem 1.2.
For the elliptic MEMS problem (1) with $N\geq 2$, $F=0$ and $\alpha>-2$, or $N\geq 2$, $F\neq 0$ and $2<\alpha<2p$, there exist infinitely many non-radial positive asymptotic radial rupture solutions satisfying
$u(x)=\Lambda|x|^{\frac{\alpha+2}{p+1}}+\sum_{j=1}^{\infty}\sum_{i=0}^{j-1}C_{ji}\left(\frac{x}{|x|}\right)(\ln|x|)^{i}|x|^{{\mu_{j}}+{\frac{\alpha+2}{p+1}}}=\Lambda|x|^{\frac{\alpha+2}{p+1}}\left(1+O\left(|x|^{\mu_{1}}\right)\right)$
as $|x|\rightarrow 0^{+}$, where $\Lambda$ is from (2), $\left\{\mu_{j}\right\}$ is a strictly increasing sequence of positive numbers diverging to $\infty$, and
${\mu_{1}}=\sigma_{1}^{(k)}~~~(\text{if}~~\delta^{(k-1)}<\alpha<\delta^{(k)},k=1,2,\dots)$
@@
$\mu_{1}=\begin{cases}{\sigma_{1}^{(k)}}&\text{if~ $\delta^{(k-1)}<\alpha<\delta^{(k)}$ and $\alpha<2p-(p+1)\sigma_{1}^{(k)}$, }k=1,2,\dots\\ \displaystyle\frac{2p-\alpha}{p+1}&\text{if ~$\delta^{(k-1)}<\alpha<\delta^{(k)}$ and $\alpha\geq 2p-(p+1)\sigma_{1}^{(k)}$, }k=1,2,\dots\end{cases}$
in the case $F\neq 0$, where
$\delta^{(k)}=\begin{cases}-2&\text{if $k=0,$}\\ \frac{-1}{2}(p+1)(N+2)+\frac{1}{2}\sqrt{(N-2)^{2}(p+1)^{2}+4(p+1)k(N-2+k)}&\text{if $k\geq 1$}\end{cases}$
and
@@
Remark.
For the case of $N=1$, there exists at least one solution near the origin satisfying
$u(x)=\Lambda^{{}^{\prime}}|x|^{\frac{\alpha+2}{p+1}}+\sum_{i=1}^{\infty}d_{i}x^{\frac{(2p-\alpha)i+(\alpha+2)}{p+1}}=\Lambda^{{}^{\prime}}|x|^{\frac{\alpha+2}{p+1}}\left(1+O\left(x^{\frac{2p-\alpha}{p+1}}\right)\right)\qquad\text{as }x\rightarrow 0^{+},$
where $\Lambda^{{}^{\prime}}=\left(\frac{\lambda}{\frac{\alpha+2}{p+1}\left(\frac{\alpha+2}{p+1}-1\right)}\right)^{\frac{1}{p+1}}$, since the arguments for solution in the case $N=1$ is completely similar to Theorem 1.1, we shall omit the case $N=1$.
To facilitate the proof of the subsequent theorems, we introduce the following change of variables. Let $t=\ln|x|$, $\theta=\dfrac{x}{|x|}$, then
$\Delta u=u_{rr}+\frac{N-1}{r}u_{r}+\frac{1}{r^{2}}\Delta_{\theta}u=\frac{\lambda|x|^{\alpha}}{u^{p}}+F.$ (3)
Let $z(t,\theta)=r^{-\frac{\alpha+2}{p+1}}u(x)-\Lambda$, where $\Lambda$ is from (2), then we obtain:
$z_{tt}+\left(N-2+2{\frac{\alpha+2}{p+1}}\right)z_{t}+(\alpha+2)(N-2+{\frac{\alpha+2}{p+1}})z+\Delta{}_{\theta}z=f(z)+Fe^{\frac{2p-\alpha}{p+1}t},$ (4)
where $f(z)=\dfrac{\lambda}{(z+\Lambda)^{p}}-\dfrac{\lambda}{\Lambda^{p}}+\dfrac{\lambda pz}{\Lambda^{p+1}}$.
We highlight the key innovations and distinguishing features of our study compared to the existing literature as follows:
(A) Extension to General Parameter Regimes. Existing literature has primarily addressed equation (1) within limited parameter regimes, such as the specific case $N=2,p=2$ studied in [8], or the case assuming $\lambda=1$ and $\alpha=0$ in [9]. We generalize these analyses to the framework of general $\alpha$. This extension is nontrivial as it introduces significant analytical challenges not present in the isotropic or restricted parameter cases.
(B) Overcoming Spectral Difficulties Involving Complex Eigenvalues. A major technical obstacle in the general $\alpha$ setting is that the associated linear operator may possess complex eigenvalues, whereas previous works mostly dealt with real spectrums. We overcome this difficulty by performing a more involved classification and establishing delicate estimates.
(C) Theoretical foundation for the design of MEMS models. While asymptotic expansions have been studied for simpler cases, our work establishes these precise characterizations within a general framework relevant to broader physical contexts. These results provide essential theoretical insights for the optimization of MEMS device structures, allowing for a more accurate description of rupture profiles.
The remainder of this paper is organized as follows. In Section 2, we derive asymptotic expansions of arbitrary order near the origin for the radial case. Section 3 extends these results non-radial case. Finally, in Section 4, we present the proofs of Theorem 1.1 and 1.2.
@@ 2 Asymptotic behavior for radial solution
Suppose that $u(x)$ is a positive radial rupture solution of (1) with $F\neq 0$. Assume it exists $\varepsilon>0$ such that
$u(r)=\Lambda r^{\frac{\alpha+2}{p+1}}\left(1+O\left(r^{\varepsilon}\right)\right)~~as~r\rightarrow 0^{+},$
where $\Lambda$ is the same as in theorem 1.1, set
$t=\ln r,z(t)=r^{-\frac{\alpha+2}{p+1}}u(r)-\Lambda,$
@@
Proof.
It follows from (4) that
@@
$f(z)=\lambda(\Lambda+z)^{-p}-\lambda\Lambda^{-p}+\lambda pz\Lambda^{-(p+1)}=O(z^{2})\quad~~as~z(t)\rightarrow 0.$
@@
Case1. $\alpha>\frac{(p+1)(N-2)}{2}(\sqrt{1+\frac{1}{p}}-1)-2,$
$\left\{\begin{aligned} \sigma_{1}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)+\frac{i}{2}\sqrt{4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)-\left(N-2+2\frac{\alpha+2}{p+1}\right)^{2}},\\ \sigma_{2}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)-\frac{i}{2}\sqrt{4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)-\left(N-2+2\frac{\alpha+2}{p+1}\right)^{2}},\end{aligned}\right$.
[…]
@@
[…]
$\left\{\begin{aligned} \sigma_{1}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)+\frac{1}{2}\sqrt{\left(N-2+2\frac{\alpha+2}{p+1}\right)^{2}-4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)}<0,\\ \sigma_{2}^{(0)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)-\frac{1}{2}\sqrt{\left(N-2+2\frac{\alpha+2}{p+1}\right)^{2}-4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)}<0,\end{aligned}\right$.
@@
$z(t)=O\left(e^{\frac{2p-\alpha}{p+1}t}\right)\qquad\text{as }t\to-\infty.$
@@
$z(t)=B_{1}e^{{\sigma_{1}^{(0)}}t}\int_{-\infty}^{t}e^{-\sigma_{1}^{(0)}s}\left[f(z(s))+Fe^{\frac{2p-\alpha}{p+1}s}\right]ds+B_{2}e^{{\sigma_{2}^{(0)}}t}\int_{-\infty}^{t}e^{-\sigma_{2}^{(0)}s}\left[f(z(s))+Fe^{\frac{2p-\alpha}{p+1}s}\right]ds$ (5)
We consider two conditions: $(i)\ 0<\varepsilon<\frac{2p-\alpha}{p+1}$, $(ii)\ \varepsilon\geq\frac{2p-\alpha}{p+1}$.
@@
we also consider two conditions $4\varepsilon\geq\frac{2p-\alpha}{p+1}$ and $4\varepsilon<\frac{2p-\alpha}{p+1}$ and use arguments similar to the above to obtain conclusions eventually.
@@
Therefore the proof of Theorem 2.1 is completed. ∎